Project

General

Profile

11082013 Minutes
Attending: Kanika, Micheal, Evan, Mat, Gavin (In robin hood garb), Dom, Nick, Susan, Jan, Jose, Siva

Evan: 2 hour window this morning, spill propagation started taking more than 20 seconds. Micheal is going to look for missing events. Looking in the FD logs the propagation time was also long. (NUmI triggers.)
Eric saw that the nova-trigger machine was inaccessible, needed reboot.
3 sec delays are confirmed resends. Larger spikes are successes with delays.

Detector - All hardware in 1 partition right now. New high gain warm and Cold settings loaded everywhere.

Gavin - Not much getting past cosmic raw
- FTS: We are going to get shafted. Old pending files are now coming through but will holod back the transfer of recent.

Michael - Why are the verticals ADCs higher and horizontals.
- Secret partition 4 (undefined data in the DB) - In contact with Jon


Slicer discussion